OpenStudy (anonymous):

(2x-3)/(x+4) I need domain,vertical asymptope, horizontal asymptope, x and y intercepts (2x-3)/(x+4) I need domain,vertical asymptope, horizontal asymptope, x and y intercepts @Mathematics

OpenStudy (mimi_x3):

Vertical Asymptote: D(x)= x+4= 0 x = -4 y-intercept: x=0 y=(0-3)/(0+4) = -3/4 x-intecept: y=0 ; 2x-3=0 =>x=3/2 To find the horizontal asymptote by finding the limit and mark it on a graph.

OpenStudy (anonymous):

Great, I was checking my answers. You're the greatest!

OpenStudy (mimi_x3):

Is my answers right ? xD

OpenStudy (anonymous):

Great, I was checking my answers. You're the greatest!

OpenStudy (anonymous):

They match mine so I hope they are. lol

OpenStudy (mimi_x3):

lols, i havent done those type of questions in a while so im not sure xD

OpenStudy (anonymous):

Great, I was checking my answers. You're the greatest!

OpenStudy (anonymous):

I am so lost in algebra its not even funny

OpenStudy (mimi_x3):

lols, algebra is easy.

OpenStudy (anonymous):

Great, I was checking my answers. You're the greatest!

OpenStudy (anonymous):

Thats what you say. I have a quiz online to do and I know I am going to bomb on it.

OpenStudy (mimi_x3):

You wil pwn it xDD If you need help I can help you.

OpenStudy (anonymous):

Great, I was checking my answers. You're the greatest!

OpenStudy (anonymous):

I would love help but it is a timed quiz and it's on my schools' site. I get 20 minutes to do 5 problems but it takes me that long to do one. lol

OpenStudy (mimi_x3):

Oh okay, thats alot of time though. Good luck anyway.

OpenStudy (anonymous):

Thanks!!!